The Royal Fruit Company produces two types of fruit drinks. The first type is 65% pure fruit juice, and the second type is 85% pure fruit juice. The company is attempting to produce a fruit drink that contains 80% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 100 pints of a mixture that is 80% pure fruit juice?

Answer by lwsshak3(11628)

The Royal Fruit Company produces two types of fruit drinks. The first type is 65% pure fruit juice, and the second type is 85% pure fruit juice. The company is attempting to produce a fruit drink that contains 80% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 100 pints of a mixture that is 80% pure fruit juice?
**
let x=pints of 65% fruit juice to be mixed
500-x=pints of 85% fruit juice to be mixed
..
65%x+85%(500-x)=80%(500)
.65x+425-.85x=400
.20x=25
x=25/.2=125
500-x=375
ans:
pints of 65% fruit juice to be mixed=125
pints of 85% fruit juice to be mixed=375